OpenStudy (anonymous):

an atom of helium-4 differs from an atom of lithium-7 in that the atom of helium-4 has 1) one more proton 2)one more neutron 3)two less protons 4)two less neutrons

OpenStudy (chmvijay):

two less neutrons!!!!!is right answer

OpenStudy (anonymous):

Helium has 2 protons and lithium has 3 protons hence no of neutrons=mass number-atomic number so,nutrons in He=4-2=2 and in Li=7-3=4. So He has two less neutrons.